# Safe Output Regulation of Coupled Hyperbolic PDE-ODE Systems

### v1

@@ 3.4 Selection of safe design parameters
where
$\acute{k}_{i}(v(t_{0}))=\frac{-1}{h_{i}(\underline{z}_{i}(\bar{t}_{0}),\bar{t}_{0})}\bigg(\sum_{\jmath=1}^{i}\frac{\partial h_{i}}{\partial z_{\jmath}}z_{\jmath+1}(\bar{t}_{0})+\frac{\partial h_{i}}{\partial t}\bigg),$ (75)
where, by (B.4) and (22), $Z(\bar{t}_{0})$ can be expressed as the initial states $z[t_{0}],w[t_{0}],Y(t_{0})$ and $v(t_{0})$. We write $\acute{k}_{i}(v(t_{0}))$ in (75) to emphasize that it depends on the unknown initial values $v(t_{0})$ of the external signal, which will be dealt with in the next section. The gain condition (74), (75) for the nonovershooting control ensures the following lemma.
@@
We give the following definition about the safe initial conation, which is the sufficient and necessary condition under which the output state stays in the safe region before the control input arrives at the distal ODE subsystem. It is the counterpart to the initial state restriction, i.e., Assumption 1, in the work [4] about safe control with delays.
@@ 3.5 Result of the state-feedback safe control
For initial data $w[t_{0}]\in L^{2}(0,1)$, $z[t_{0}]\in L^{2}(0,1)$, $Y(t_{0})\in\mathbb{R}^{n}$, for design parameters $\kappa_{i}$, $i=1,\ldots,n-1$ satisfying (74), the closed-loop system including the plant (1)–(5) with the controller (72) has the properties:
1) There exists a unique solution $w[t]\in L^{2}(0,1)$, $z[t]\in L^{2}(0,1)$, $Y(t)\in\mathbb{R}^{n}$ to the system (1)–(5) under the control input (72).
2) Tracking error $e(t)=y_{1}(t)-r(t)$ is convergent to zero and all states are bounded;
3) Safety is ensured in the sense that:
@@
c) When the safe initial condition in Definition 1 is not satisfied, and $h(e(\bar{t}_{0}),\bar{t}_{0})\leq 0$, the sate will return and stay in the safe region no latter than a finite time $\bar{t}_{0}+t_{a}$ where a constant $t_{a}>0$ can be arbitrarily assigned by users, i.e., $h(e(t),t)\geq 0,~\forall t\geq\bar{t}_{0}+t_{a}$.
@@
1) For the given initial conditions, recalling Assumption 2 and the transformations (22), (47)–(49), (62), (63), we have $\beta[t_{0}]\in L^{2}(0,1)$, $\alpha[t_{0}]\in L^{2}(0,1)$, $H(t_{0})\in\mathbb{R}^{n}$. Considering the target system (67)–(71), it is followed that there exists a unique solution $\beta[t]\in L^{2}(0,1)$, $\alpha[t]\in L^{2}(0,1)$, $H(t)\in\mathbb{R}^{n}$. According to (48), we have $\frac{\partial h_{i}}{\partial z_{i-1}}=\frac{\partial h_{i-1}}{\partial z_{i-2}}+\frac{\partial h_{i-1}}{\partial z_{i-1}\partial t}+k_{i-1}\frac{\partial h_{i-1}}{\partial z_{i-1}}=\frac{\partial h_{i-1}}{\partial z_{i-2}}+k_{i-1}\frac{\partial h_{1}}{\partial z_{1}}+\frac{\partial h_{1}}{\partial z_{1}\partial t}=\frac{\partial h_{2}}{\partial z_{1}}+\frac{\partial h_{1}}{\partial z_{1}}\sum_{j=2}^{j=i-1}k_{j}+(i-2)\frac{\partial h_{1}}{\partial z_{1}\partial t}=\frac{\partial h_{1}}{\partial z_{1}}\sum_{j=1}^{j=i-1}k_{j}+(i-1)\frac{\partial h_{1}}{\partial z_{1}\partial t},~~\forall i\geq 2.$ Recursively, we obtain $\sum_{i=1}^{n}\sum_{j=1}^{i}|\frac{\partial h_{i}}{\partial z_{j}}|\leq\mathcal{M}\big(|\frac{\partial h_{1}}{\partial z_{1}}|+\sum_{j=1}^{n-1}|\frac{\partial h_{1}}{\partial z_{1}\partial t^{n-1}}|\big)$ for some positive constant $\mathcal{M}$. Recalling Assumption 2, (47), and (49) that renders that $\sigma$ has continuous derivatives of all orders, we know $\frac{\partial h_{i}}{\partial z_{j}}$, $1\leq j\leq i\leq n$, $i,j\in\mathbb{N}$ are bounded. According to (47), (49), and Assumption 2, recursively applying (48), we obtain from $H(t)\in\mathbb{R}^{n}$ that $Z(t)\in\mathbb{R}^{n}$. It is then obtained that $Y(t)\in\mathbb{R}^{n}$, recalling (22). According to the theory of Volteral integral equations [32], considering the backstepping transformation (62), (63) as well as the boundedness and continuity of the kernels given in Appendix 0.3, there exists a bounded and continuous the kernel $\mathcal{K}(x,y)$ such that
@@
$+\mathcal{K}_{1}(x)Y(t)+\mathcal{K}_{2}(x)v(t)+\left(\begin{array}[]{c}q(x,t)\\ p(x,t)\\ \end{array}\right)$ (85)
$+\int_{0}^{x}\mathcal{K}(x,y)\left(\begin{array}[]{c}q(y,t)\\ p(y,t)\\ \end{array}\right)dy$ (88)
where $\mathcal{K}(x,y)$ can be computed from the kernels ${\phi},\varphi,\Psi,\Phi$, and where $\mathcal{K}_{1}(x)=\left(\gamma(x),\lambda(x)\right)^{T}+\int_{0}^{x}\mathcal{K}(x,y)\left(\gamma(y),\lambda(y)\right)^{T}dy$, $\mathcal{K}_{2}(x)=\left(\bar{\gamma}(x),\bar{\lambda}(x)\right)^{T}+\int_{0}^{x}\mathcal{K}(x,y)\left(\bar{\gamma}(y),\bar{\lambda}(y)\right)^{T}dy$. According to the exogenous signal defined in Sec. 2.2, we have $v(t)\in\mathbb{R}^{n_{v}}$. Applying (88), together with (64), (65), Property 1 is obtained.
2) Recalling the target system (67)–(71) and Assumption 2 indicates that $\vartheta$ is bounded, by the method of characteristics, it is obtained that $\beta[t],\alpha[t],H(t)$ are bounded all the time, and moreover, $\beta[t]\equiv 0$, $\alpha[t]\equiv 0$ for $t\geq\frac{1}{q_{1}}+\frac{1}{q_{2}}$ and $|H(t)|$ is exponentially convergent to zero recalling the fact that $A_{\rm h}$ defined in (61) is Hurwitz. We then obtain from (47), (49), and Assumption 2 that $z_{1}$ is bounded. Recursively applying (48), recalling the fact that $\frac{\partial h_{i}}{\partial z_{j}}$, $1\leq j\leq i\leq n$, $i,j\in\mathbb{N}$ are bounded, which is obtained in the proof of property 1, we obtain $|Z(t)|$ is bounded. Recalling (22) and the boundedness of $|v(t)|$ straightforwardly obtained for Sec. 2.2, we have that $|Y(t)|$ is bounded. Applying the Cauchy-Schwarz inequality into the inverse transformation (88), we have that $z[t],w[t]$ are bounded. Besides, it is obtained from the exponential convergence to zero of $h_{1}$ that $h(e,t)$ is exponentially convergent to zero according to (47) and (49) that shows $\sigma(t)\equiv 0$ after a finite time $t=\bar{t}_{0}+t_{a}$. Recalling Assumption 2, we know $e(t)$ is exponentially convergent to zero. Property 2 is obtained.
3) According to the choice of the design parameters $\kappa_{i}$, $i=1,\ldots,n-1$ in (74), recalling Lemma 1, we know $h_{i}>0$ at $t=\bar{t}_{0}=\frac{1}{q_{2}}$. Because $\beta(0,t)\equiv 0$ for $t\geq\frac{1}{q_{2}}$ considering the target system (67)–(71), and the system matrix $A_{\rm h}$ defined in (61) represents a form of high-relative-degree CBFs $h_{i}$ as presented in [15], [18], we have $h_{i}>0$ for $t\geq\bar{t}_{0}$. Based on this, we prove the safety in cases a)–c) as follows:
@@
c) When the safe initial conditions in Definition 1 is not satisfied, if $h(e(\bar{t}_{0}),\bar{t}_{0})\leq 0$, it is followed that $\sigma(t)\equiv 0$ after a finite time $t=\bar{t}_{0}+t_{a}$ according to (49) where $t_{a}>0$ is a free design parameter. We then obtain from the fact $h_{1}>0$ for $t\geq\bar{t}_{0}$ and (47) that $h(e(t),t)\geq 0$ for $t\geq\bar{t}_{0}+t_{a}$. Property 3 is thus obtained. The proof of this theorem is complete.
4 Unmeasured Distributed States and Unknown External Disturbances
@@ 4.1 State observer and disturbance estimator
In the last section, we have proposed the nominal control design on the basis of a completely known model. In practice, the distributed states in PDE, and also the external disturbances, are always inaccessible. Therefore, we use the measurements at the boundaries, i.e., $Y(t)$, $z(1,t)$, together with the known reference signal $r(t)$, to build the state and disturbance observer to estimate the unmeasured distributed states and the unknown external disturbance:
@@
with the initial states $\hat{z}[t_{0}],\hat{w}[t_{0}],\hat{v}(t_{0})$ chosen in the known bounds given in Assumptions 3, 4, where $\hat{v}_{r},\hat{v}_{d}]$ is the estimate of $v^{T}=[v_{r},v_{d}]$, and $\hat{z}$, $\hat{w}$ are observer states. $\bar{G}_{i}=G_{i}\bar{P}_{d}$, $i=2,3,4,5$. The vector $L_{r}$ is selected such that $S_{r}-L_{r}\bar{P}_{r}$ Hurwitz, considering that $(S_{r},\bar{P}_{r})$ is observable, and observer gains $L_{d}$, $L_{1}(x)$, $L_{2}(x)$ are defined later. Define the estimation error states
@@
$\tilde{z}(0,t)$ $=p\tilde{w}(0,t)+\bar{G}_{4}\tilde{v}_{d}(t),$ (100)
@@
recalling (1)–(5), (13)–(15). Applying the transformation:
@@
where $K^{11}(x,y)$, $K^{12}(x,y)$, $K^{21}(x,y)$, $K^{22}(x,y)$ on $\{(x,y)|0\leq x\leq y\leq 1\}$ satisfy $q_{1}K^{11}_{x}(x,y)+q_{1}K^{11}_{y}(x,y)=d_{1}K^{21}(x,y)$, $q_{1}K^{12}_{x}(x,y)-q_{2}K^{12}_{y}(x,y)=d_{1}K^{22}(x,y)$, $q_{2}K^{21}_{x}(x,y)-q_{1}K^{21}_{y}(x,y)=-d_{2}K^{11}(x,y)$, $q_{2}K^{22}_{x}(x,y)+q_{2}K^{22}_{y}(x,y)=-d_{2}K^{12}(x,y)$, $K^{21}(x,x)=\frac{-d_{2}}{q_{1}+q_{2}}$, $K^{12}(x,x)=\frac{d_{1}}{q_{1}+q_{2}}$, $K^{11}(0,y)=pK^{21}(0,y)$, $K^{22}(0,y)=\frac{1}{p}K^{12}(0,y)$, whose solution is in the same form as (C.18), (C.19) (which can be seen clearly by changing the domain from $\{(x,y)|0\leq x\leq y\leq 1\}$ to $\{(x,y)|0\leq y\leq x\leq 1\}$ and comparing it with (B.4)–(B.6) of [24]), and choosing the observer gains as
@@
$\tilde{\alpha}(0,t)$ $=p\tilde{\beta}(0,t)+\bar{G}_{4}\tilde{v}_{d}(t),$ (110)
where $\bar{K}_{1}(x),\bar{K}_{2}(x)$ satisfy $\bar{K}_{1}(x)-\int_{x}^{1}K^{22}(x,y)\bar{K}_{1}(y)dy-\int_{x}^{1}K^{21}(x,y)\bar{K}_{2}(y)dy-q_{2}K^{22}(x,1)\bar{G}_{5}-\bar{G}_{3}(x)=0$, and $\bar{K}_{2}(x)-\int_{x}^{1}K^{11}(x,y)\bar{K}_{2}(y)dy-\int_{x}^{1}K^{12}(x,y)\bar{K}_{1}(y)dy-q_{2}K^{12}(x,1)\bar{G}_{5}-\bar{G}_{2}(x)=0$.

### v2

@@ 3.4 Selection of safe design parameters
where $\acute{k}_{i}(v(t_{0}))=\frac{-1}{h_{i}(\underline{z}_{i}(\bar{t}_{0}),\bar{t}_{0})}\big(\sum_{\jmath=1}^{i}\frac{\partial h_{i}}{\partial z_{\jmath}}z_{\jmath+1}(\bar{t}_{0})+\frac{\partial h_{i}}{\partial t}\big)$. By (B.5) and (22), $Z(\bar{t}_{0})$ can be expressed as the initial states $z[t_{0}],w[t_{0}],Y(t_{0})$ and $v(t_{0})$. We write $\acute{k}_{i}(v(t_{0}))$ to emphasize that it depends on the unknown initial values $v(t_{0})$ of the external signal, which will be dealt with in the next section. The gain condition (73) for the nonovershooting control ensures the following lemma.
@@
We give the following definition about the safe initial condition, which is the sufficient and necessary condition under which the output state stays in the safe region before the control input arrives at the distal ODE subsystem. It is the counterpart to the initial state restriction, i.e., Assumption 1, in the work [4] about safe control with delays.
@@ 3.5 Result of the state-feedback safe control
Assume that the closed-loop system consisting of (1)–(5) and (71) admits a classical solution. For initial data $w[t_{0}]\in L^{\infty}(0,1)$, $z[t_{0}]\in L^{\infty}(0,1)$, $Y(t_{0})\in\mathbb{R}^{n}$, and for design parameters $\kappa_{i}$, $i=1,\ldots,n-1$, satisfying (73), the following properties hold:
1) Tracking error $e(t)=y_{1}(t)-r(t)$ is convergent to zero and all states are bounded;
2) Safety is ensured in the sense that:
@@
c) When the safe initial condition in Definition 1 is not satisfied, and $h(e(\bar{t}_{0}),\bar{t}_{0})\leq 0$, the state will return and stay in the safe region no latter than a finite time $\bar{t}_{0}+t_{a}$ where a constant $t_{a}>0$ can be arbitrarily assigned by users, i.e., $h(e(t),t)\geq 0,~\forall t\geq\bar{t}_{0}+t_{a}$.
@@
1) Recalling the target system (68)–(70) and Assumption 3 indicates that $\vartheta$ is bounded, by the method of characteristics, it is obtained that $\beta[t],\alpha[t],H(t)$ are bounded all the time, and moreover, $\beta[t]\equiv 0$, $\alpha[t]\equiv 0$ for $t\geq\frac{1}{q_{1}}+\frac{1}{q_{2}}$ and $|H(t)|$ is exponentially convergent to zero recalling the fact that $A_{\rm h}$ defined in (62) is Hurwitz.
We then obtain from (47), (49), and Assumption 3 that $z_{1}$ is convergent to zero. According to (48), we have $\frac{\partial h_{i}}{\partial z_{i-1}}=\frac{\partial h_{i-1}}{\partial z_{i-2}}+\frac{\partial h_{i-1}^{2}}{\partial z_{i-1}\partial t}+k_{i-1}\frac{\partial h_{i-1}}{\partial z_{i-1}}=\frac{\partial h_{i-1}}{\partial z_{i-2}}+k_{i-1}\frac{\partial h_{1}}{\partial z_{1}}+\frac{\partial h_{1}^{2}}{\partial z_{1}\partial t}=\frac{\partial h_{2}}{\partial z_{1}}+\frac{\partial h_{1}}{\partial z_{1}}\sum_{j=2}^{j=i-1}k_{j}+(i-2)\frac{\partial h_{1}^{2}}{\partial z_{1}\partial t}=\frac{\partial h_{1}}{\partial z_{1}}\sum_{j=1}^{j=i-1}k_{j}+(i-1)\frac{\partial h_{1}^{2}}{\partial z_{1}\partial t},~~\forall i\geq 2,$ where (53) has been recalled. Repeating the same process for $\frac{\partial h_{i}}{\partial z_{i-j}}$, $j+1\leq i\leq n$, $j=2,\cdots,n-1$, we obtain $\sum_{i=1}^{n}\sum_{j=1}^{i}|\frac{\partial h_{i}}{\partial z_{j}}|\leq\mathcal{M}\big(|\frac{\partial h_{1}}{\partial z_{1}}|+\sum_{j=1}^{n-1}|\frac{\partial h_{1}^{n}}{\partial z_{1}\partial t^{n-1}}|\big)$ for some positive constant $\mathcal{M}$. Recalling Assumption 3, (47), and (49) that renders that $\sigma$ has continuous derivatives of all orders, it means that $\frac{\partial h_{i}}{\partial z_{j}}$, $1\leq j\leq i\leq n$, $i,j\in\mathbb{N}$ are bounded. Applying (48) for $i=2$, recalling (53), we have $z_{2}=\frac{1}{\vartheta}({h_{2}-\frac{\partial h_{1}}{\partial t}-k_{1}h_{1}})$, which is bounded according to the boundedness of $h_{1},h_{2}$, and $\frac{\partial h_{1}}{\partial t}$ which is ensured by the fact that $h_{1}$ is $n$ times differentiable in Assumption 3. Applying (48) for $i=3$, it is obtained that $z_{3}=\frac{1}{\vartheta}({h_{3}-\frac{\partial h_{2}}{\partial t}-k_{2}h_{2}-\frac{\partial h_{2}}{\partial z_{1}}z_{2}})$ is bounded, where $\frac{\partial h_{2}}{\partial t}=\frac{\partial h_{1}^{2}}{\partial z_{1}\partial t}z_{2}+\frac{\partial h_{1}^{2}}{\partial t^{2}}+k_{1}\frac{\partial h_{1}}{\partial t}$ is bounded recalling that $z_{2}$ is bounded as proved above as well as Assumption 3. Similarly, applying (48) for $i=4$, we have that $z_{4}=\frac{1}{\vartheta}({h_{4}-\frac{\partial h_{3}}{\partial t}-k_{3}h_{3}-\frac{\partial h_{3}}{\partial z_{1}}z_{2}-\frac{\partial h_{3}}{\partial z_{2}}z_{3}})$ is bounded, where $\frac{\partial h_{3}}{\partial t}$, which can be expressed by $\frac{\partial h_{1}^{i}}{\partial t^{i}}$ and $z_{i}$, $i=1,2,3$, is bounded according to Assumption 3 and the boundedness of $z_{i}$, $i=1,2,3$ proved above. Recursively applying (48), and repeating the above process until $z_{n}$, we have the inverse transformation of (48) from $H$ to $Z$, and obtain that $|Z(t)|$ is bounded for the boundedness of $|H(t)|$. Recalling (22) and the boundedness of $|v(t)|$ straightforwardly obtained for Sec. 2.2, we have that $|Y(t)|$ is bounded.
According to the theory of Volteral integral equations [38], considering the backstepping transformation (63), (64) as well as the boundedness and continuity of the kernels given in Appendix C, there exists a bounded and continuous the kernel $\mathcal{K}(x,y)$ such that
@@
$+\mathcal{K}_{1}(x)Y(t)+\mathcal{K}_{2}(x)v(t)+\left(\begin{array}[]{c}\bar{\varsigma}(x,t)\\ {\varsigma}(x,t)\\ \end{array}\right)$ (83)
$+\int_{0}^{x}\mathcal{K}(x,y)\left(\begin{array}[]{c}\bar{\varsigma}(y,t)\\ {\varsigma}(y,t)\\ \end{array}\right)dy$ (86)
where $\mathcal{K}(x,y)$ can be computed from the kernels ${\phi},\varphi,\Psi,\Phi$, and where $\mathcal{K}_{1}(x)=\left(\gamma(x),\lambda(x)\right)^{T}+\int_{0}^{x}\mathcal{K}(x,y)\left(\gamma(y),\lambda(y)\right)^{T}dy$, $\mathcal{K}_{2}(x)=\left(\bar{\gamma}(x),\bar{\lambda}(x)\right)^{T}+\int_{0}^{x}\mathcal{K}(x,y)\left(\bar{\gamma}(y),\bar{\lambda}(y)\right)^{T}dy$. Applying the Cauchy-Schwarz inequality to the inverse transformation (86), we have that $z[t],w[t]$ are bounded. Besides, it is obtained from the exponential convergence to zero of $h_{1}$ that $h(e,t)$ is exponentially convergent to zero according to (47) and (49) that shows $\sigma(t)\equiv 0$ after a finite time $t=\bar{t}_{0}+t_{a}$. Recalling Assumption 3, we know $e(t)$ is exponentially convergent to zero. Property 1 is obtained.
2) According to the choice of the design parameters $\kappa_{i}$, $i=1,\ldots,n-1$ in (73), recalling Lemma 1, we know $h_{i}>0$ at $t=\bar{t}_{0}=\frac{1}{q_{2}}$. Because $\beta(0,t)\equiv 0$ for $t\geq\frac{1}{q_{2}}$ considering the target system (68)–(70), and the system matrix $A_{\rm h}$ defined in (62) represents a form of high-relative-degree CBFs $h_{i}$ as presented in [15], [22], we have $h_{i}>0$ for $t\geq\bar{t}_{0}$. Based on this, we prove the safety in cases a)–c) as follows:
@@
c) When the safe initial conditions in Definition 1 is not satisfied, if $h(e(\bar{t}_{0}),\bar{t}_{0})\leq 0$, it is followed that $\sigma(t)\equiv 0$ after a finite time $t=\bar{t}_{0}+t_{a}$ according to (49) where $t_{a}>0$ is a free design parameter. We then obtain from the fact $h_{1}>0$ for $t\geq\bar{t}_{0}$ and (47) that $h(e(t),t)\geq 0$ for $t\geq\bar{t}_{0}+t_{a}$. Property 2 is thus obtained. The proof of this theorem is complete.
Remark 3.
If the reference $r(t)\equiv 0$, and the chosen barrier function $h$ satisfying that $\frac{\partial h^{i}(e(t),t)}{\partial t^{i}}$, $i=1,\cdots,n$, converge to zero once $e(t)$ converges to zero, our controller can achieve that full states in the closed-loop system, including the PDE and ODE, are convergent to zero, meanwhile, the output state $C_{1}Y$ stays in the safe region.
4 Unmeasured States and Unknown Disturbances
@@ 4.1 State observer and disturbance estimator
In the last section, we have proposed the nominal control design on the basis of a completely known model. In practice, the full states, and also the external disturbances, are always inaccessible. Therefore, we use the measurements $y_{1}(t)=C_{1}Y(t)$, $z(1,t)$, together with the known reference signal $r(t)$, to build an extended observer to estimate the unmeasured states and the unknown external disturbance:
@@
with the initial states $\hat{z}[t_{0}],\hat{w}[t_{0}],\hat{v}(t_{0})$ chosen in the known bounds given in Assumptions 4, 5, where $\hat{v}_{r},\hat{v}_{d}]$ is the estimate of external signals $v^{T}=[v_{r},v_{d}]$, and $\hat{z}$, $\hat{w}$, $\hat{Y}$ are observer states. The gains $\bar{G}_{i}=G_{i}\bar{P}_{d}$, $i=2,3,4,5$. The observer gain $L_{y}$ is selected such that $A-L_{y}C_{1}$ Hurwitz, considering that $(A,C_{1})$ is observable in Assumption 2. Ohter observer gains $L_{r}$, $L_{y}$, $L_{d}$, $L_{1}(x)$, $L_{2}(x)$ are defined later. Define the estimation error states
@@
$\tilde{z}(0,t)$ $=p\tilde{w}(0,t)+C\tilde{Y}(t)+\bar{G}_{4}\tilde{v}_{d}(t),$ (99)
@@
recalling (1)–(5), (13)–(15). Introduce the transformation:
@@
${\tilde{Y}(t)=}$ ${\tilde{X}(t)-\int_{0}^{1}K_{0}(x)\tilde{\alpha}(x,t)dx-\int_{0}^{1}K_{1}(x)\tilde{\beta}(x,t)dx}$ (103)
where $K^{11}(x,y)$, $K^{12}(x,y)$, $K^{21}(x,y)$, $K^{22}(x,y)$ on $\{(x,y)|0\leq x\leq y\leq 1\}$ satisfy $q_{1}K^{11}_{x}(x,y)+q_{1}K^{11}_{y}(x,y)=d_{1}K^{21}(x,y)$, $q_{1}K^{12}_{x}(x,y)-q_{2}K^{12}_{y}(x,y)=d_{1}K^{22}(x,y)$, $q_{2}K^{21}_{x}(x,y)-q_{1}K^{21}_{y}(x,y)=-d_{2}K^{11}(x,y)$, $q_{2}K^{22}_{x}(x,y)+q_{2}K^{22}_{y}(x,y)=-d_{2}K^{12}(x,y)$, $K^{21}(x,x)=\frac{-d_{2}}{q_{1}+q_{2}}$, $K^{12}(x,x)=\frac{d_{1}}{q_{1}+q_{2}}$, $K^{11}(0,y)=pK^{21}(0,y)-CK_{0}(y)$, $K^{22}(0,y)=\frac{1}{p}K^{12}(0,y)-\frac{1}{p}CK_{1}(y)$, with $BK^{21}(0,x)-K_{0}^{\prime}(x)q_{1}+(A-L_{y}C_{1})K_{0}(x)=0$, $BK^{22}(0,x)+K_{1}^{\prime}(x)q_{2}+(A-L_{y}C_{1})K_{1}(x)=0$, $B-K_{1}(0)q_{2}=0$, $K_{0}(0)=0$, whose solution can be obtained as the same process in Appendix C (which can be seen clearly by changing the domain from $\{(x,y)|0\leq x\leq y\leq 1\}$ to $\{(x,y)|0\leq y\leq x\leq 1\}$ and comparing it with (B.4)–(B.6) of [30]). Applying the transformations (101)–(103) and choosing the observer gains as
@@
$\tilde{\alpha}(0,t)$ $=p\tilde{\beta}(0,t)+C\tilde{X}(t)+\bar{G}_{4}\tilde{v}_{d}(t),$ (111)
where $\bar{P}_{1}=\int_{0}^{1}K_{0}(x)p_{1}(x)dx+\int_{0}^{1}K_{1}(x)p_{2}(x)dx-K_{0}(1)q_{1}$, and where $\bar{K}_{1}(x),\bar{K}_{2}(x)$ satisfy $\bar{K}_{1}(x)-\int_{x}^{1}K^{22}(x,y)\bar{K}_{1}(y)dy$ $\int_{x}^{1}K^{21}(x,y)\bar{K}_{2}(y)dy-q_{2}K^{22}(x,1)\bar{G}_{5}-\bar{G}_{3}(x)=0$, and $\bar{K}_{2}(x)-\int_{x}^{1}K^{11}(x,y)\bar{K}_{2}(y)dy-\int_{x}^{1}K^{12}(x,y)\bar{K}_{1}(y)dy-q_{2}K^{12}(x,1)\bar{G}_{5}-\bar{G}_{2}(x)=0$, and where $\bar{K}_{0}=\int_{0}^{1}K_{0}(x)\bar{K}_{2}(x)dx$ $\bar{G}_{1}+\int_{0}^{1}K_{1}(x)\bar{K}_{1}(x)dx+K_{1}(1)q_{2}\bar{G}_{5}$.